OpenStudy (anonymous):

an 82 inch length of ribbon is to be cut into three pieces. The longest piece is to be 38 inches longer than the shorter piece, and the third piece is to be half the length of the longest piece. Find the length of each piece of ribbon.

OpenStudy (anonymous):

to set this up would it be like 82=38+s1/2?

hartnn (hartnn):

let longest piece be s. shortest piece = s-38 third piece = s/8. sum = 82 so s+s-38+s/8 = 82 can u solve this for s ?

OpenStudy (anonymous):

um let me try

hartnn (hartnn):

sorry! typo

hartnn (hartnn):

third piece is half --->s/2 s+s-38+s/2 = 82

OpenStudy (anonymous):

s=48? so let me see

hartnn (hartnn):

yup, s=48. good work :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

wait i got 58+24=82 did i miss something?

hartnn (hartnn):

longest piece = s =48 smallest piece = s-38 = 10 third = s/2 = 24 48+10+24 =82 thats correct........

OpenStudy (anonymous):

o.O idk where i went wrong then

OpenStudy (anonymous):

oh i see i got s=48 but didn't write it properly and look at the signs xD

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome again :)